OpenStudy (anonymous):

Write the equation of each line in slope-intercept form. x=-2, 1, 4 , 7 f(x)= -14, -5, 4, 13

OpenStudy (anonymous):

These numbers were shown in a box, one over the other. for example, -2 is in the x row over -14 in the f(x) row

myininaya (myininaya):

do you know how to find the slope?

OpenStudy (anonymous):

I'm just confused on where to begin

myininaya (myininaya):

\[m=\frac{y_2-y_1}{x_2-x_1}\] this is formula for slope

myininaya (myininaya):

\[m=\frac{-14-(-5)}{-2-1}\]

myininaya (myininaya):

the bottom row is the y's top row is the x's

myininaya (myininaya):

sorry i was having some internet issues

myininaya (myininaya):

\[m=\frac{-14+5}{-3}=\frac{-9}{-3}=3\]

myininaya (myininaya):

all i did was simplify the slope i had above

myininaya (myininaya):

now you need the y-intercept

myininaya (myininaya):

so first of all slope intercept form looks like this \[y-y_1=m(x-x_1)\]

myininaya (myininaya):

we know m from above

myininaya (myininaya):

so we have \[y-y_1=3(x-x_1)\]

myininaya (myininaya):

now enter any of the points you were given and you are done you could enter (-2,-14) or (1,-5) or (4,4) or (7,13)